I need two different ways to solve this equation . -4x-2=4
Explain which method you prefer and why

Answer by stanbon(75887)

-4x-2=4
-----
-4x = 4+2
-4x = 6
x = 6/-4 =-3/2
------------
-4x-2 = 4
Divide thru by -2 to get:
2x +1 = -2
2x = -3
x = -3/2
============
Cheers,
Stan H.
============